Calculus II, Spring 2013 (http://www.math.nthu.edu.tw/˜ wangwc/)

Homework Assignment for Week 09

1. Section 13.6: Problem 15 (Read Example 5 for the meaning of direction of A), 21, 23, 25, 27, 37, 39, 47, 49, 51, 53, 54, 57, 58.

2. (s13.5 extra3) Find the Taylor expansion of f(x, y, z) around (x_{0}, y_{0}, z_{0}) up to quadratic terms of x, y and z. Give an expression of the remainder term, \frac{1}{3!}(⋯).